OpenStudy (anonymous):

Use the technique of completing the square to transform the quadratic equation into the form (x + c)2 = a. 2x2 – 28x + 84 = 0

OpenStudy (paxpolaris):

\[\large 2x^2 – 28x + 84 = 0\] move 84 to the other side\[\implies \large 2x^2 – 28x =- 84\] divide everything by 2\[\implies \large x^2 – 14x =- 42\]

OpenStudy (paxpolaris):

next you complet the square

OpenStudy (paxpolaris):

\[\implies \Large x^2 – 14x +49=- 42+49\]

OpenStudy (paxpolaris):

\[\implies \Large \left( x-7 \right)^2=7\]

OpenStudy (anonymous):

Thank you, so much.